3. Lightfoot Inc., a software development firm, has stock outstanding as follows: 30,000 shares of cumulative preferred 2% stock, $20 par, and 38,000 shares of $125 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $4,500; second year, $7,500; third year, $54,020; fourth year, $114,220.Calculate the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) $ $ $ $
Common stock (dividend per share) $ $ $ $
Answer #1

Cumulative preferrence share are those share on which any outstanding dividend of previous year is to be paid in the next year before any dividend is distributed to the common share holders.

Preferrence share = 30,000*20 = $600,000

Preferrence dividend = 600,000*2% = $12,000

First year

Total dividend paid = $4,500

Preferrence share holder dividend = $4,500

Preferrence share holder dividend per share = 4,500/30,000 = $0.15

Common share holder dividend = 0

Outstanding preference shares holder dividend = 12,000 - 4,500 = $7,500

Second year

Total dividend paid = $7,500

Preferrence share holder dividend =$7,500

Preferrence share holder dividend per share = 7,500/30,000 = $0.25

Common share holder dividend = 0

Outstanding preference shares holder dividend = 12,000 - 7,500 = 4,500 and previous outstanding of $7,500 total = $12,000

Third year

Total dividend paid = $54,020

Preferrence share holder dividend = 12,000 + 12,000 = $24,000 ( Previous outstanding $12,000 and this year 12,000)

Preferrence share holder dividend per share = 24,000/30,000 = $0.80

Common share holder dividend =54,020 - 24,000 = $30,020

Common share holders dividend per share = 30,020/38,000 = $0.79

Fourth year

Total dividend paid = 114,220

Preferrence share holder dividend = 12,000

Preferrence share holder dividend per share = 12,000/30,000 = $0.40

Common share holder dividend =114,220 - 12,000 = $102,220

Common share holders dividend per share = 102,220/38,000 = $2.69

If you find the answer helpful please upvote.